OpenStudy (chris215):

A ball is thrown vertically upward from the top of a 100 foot tower, with an initial velocity of 10 ft/sec. Its position function is s(t) = -16t2 + 10t + 100. What is its velocity in ft/sec when t = 2 seconds?

OpenStudy (superdavesuper):

getting sick of seeing me yet, chris? :P Velocity = d(position function)/dt = ds/dt = d(-16t2 + 10t + 100)/dt = ....

OpenStudy (chris215):

-32t+10 -54

OpenStudy (misty1212):

looks good to me!

OpenStudy (chris215):

thanks guys!

OpenStudy (superdavesuper):

@chris215 GREAT job - u did most of the work here :)